# Discovering Mixtures of Structural Causal Models from Time Series Data

### v2

@@ D.5 Aggregating the temporal adjacency matrix across time
The Netsim and DREAM3 datasets used in the evaluation provide ground-truth time-aggregated causal graphs. In order to make our model output comparable, we follow the procedure outlined in [Gong et al., 2022] to convert the time-lag adjacency matrix to an aggregated matrix. The $(i,j)^{\text{th}}$ entry of the aggregated matrix $\mathcal{G}_{\text{agg}}$ is 1 iff $\mathcal{G}_{\ell}^{ij}=1$ for some lag value $\ell$ in the time-lag matrix $\mathcal{G}$. Both Rhino and MCD represent the edges as Bernoulli random variables, hence output a probability score for each edge. For evaluating the F1 score, we threshold the probability values at 0.5, i.e., edges with a probability $\geq$ 0.5 are considered as predicted edges.
@@ Appendix E Toy example
However, if the graph membership of the samples is unknown, inferring a single causal graph to explain the causal relationships from the dataset would result in spurious causal relationships. For example, going by conditional independence tests, note that none of the nodes would be conditionally independent of each other for any conditioning set. This is also shown in the output of the ${}^{+}$ algorithm, where a dense graph is inferred as shown in Figure 21. Thus, it is crucial to use a mixture distribution to model observational data from such heterogeneous distributions.
Figure 21: ${}^{+}$ output on the toy-example. The algorithm infers a dense graph with many spurious causal relationships.

### v3

@@ D.5 Aggregating the temporal adjacency matrix across time
The Netsim and DREAM3 datasets used in the evaluation provide ground-truth time-aggregated causal graphs. In order to make our model output comparable, we follow the procedure outlined in (Gong et al., 2022) to convert the time-lag adjacency matrix to an aggregated matrix. The $(i,j)^{\text{th}}$ entry of the aggregated matrix $\mathcal{G}_{\text{agg}}$ is 1 iff $\mathcal{G}_{\ell}^{ij}=1$ for some lag value $\ell$ in the time-lag matrix $\mathcal{G}$. Both Rhino and MCD represent the edges as Bernoulli random variables, hence output a probability score for each edge. For evaluating the F1 score, we threshold the probability values at 0.5, i.e., edges with a probability $\geq$ 0.5 are considered as predicted edges.
@@ D.6 Pair-wise graph distance in the mixture distributions
Synthetic-Perturbed $(p=0.005)$ 10 5 2.40 0.92 1 4
Synthetic-Perturbed $(p=0.008)$ 10 5 5.20 0.98 4 7
Synthetic-Perturbed $(p=0.01)$ 10 5 7.00 0.89 5 8
Synthetic-Perturbed $(p=0.05)$ 10 5 21.40 5.37 15 30
Synthetic-Perturbed $(p=0.1)$ 10 5 48.20 2.64 45 53
Synthetic-Perturbed $(p=0.005)$ 10 10 2.44 1.04 1 5
Synthetic-Perturbed $(p=0.008)$ 10 10 3.56 1.36 1 6
Synthetic-Perturbed $(p=0.01)$ 10 10 6.11 2.51 1 12
Synthetic-Perturbed $(p=0.05)$ 10 10 23.04 4.17 15 32
Synthetic-Perturbed $(p=0.1)$ 10 10 47.56 8.56 29 65
@@ Appendix E Toy example
However, if the graph membership of the samples is unknown, inferring a single causal graph to explain the causal relationships from the dataset would result in spurious causal relationships. For example, going by conditional independence tests, note that none of the nodes would be conditionally independent of each other for any conditioning set. This is also shown in the output of the PCMCI+ algorithm, where a dense graph is inferred as shown in Figure 22. Thus, it is crucial to use a mixture distribution to model observational data from such heterogeneous distributions.
Figure 22: PCMCI+ output on the toy-example. The algorithm infers a dense graph with many spurious causal relationships.